# Name Tagging Under Domain Shift via Metric Learning for Life Sciences

### v1

@@ Abstract
Name tagging is a key component of Information Extraction (IE), particularly in scientific domains such as biomedicine and chemistry, where large language models (LLMs), e.g., ChatGPT, fall short. We investigate the applicability of transfer learning for enhancing a name tagging model trained in the biomedical domain (the source domain) to be used in the chemical domain (the target domain). A common practice for training such a model in a few-shot learning setting is to pretrain the model on the labeled source data, and then, to finetune it on a hand-full of labeled target examples. In our experiments we observed that such a model is prone to mis-labeling the source entities, which can often appear in the text, as the target entities. To alleviate this problem, we propose a model to transfer the knowledge from the source domain to the target domain, however, at the same time, to project the source entities and target entities into separate regions of the feature space. This diminishes the risk of mis-labeling the source entities as the target entities. Our model consists of two stages: 1) entity grouping in the source domain, which incorporates knowledge from annotated events to establish relations between entities, and 2) entity discrimination in the target domain, which relies on pseudo labeling and contrastive learning to enhance discrimination between the entities in the two domains. We carry out our extensive experiments across three source and three target datasets, and demonstrate that our method outperforms the baselines, in some scenarios by 5% absolute value.
Name Tagging Under Domain Shift via Metric Learning
@@
[…]
@@ 1 Introduction
Name tagging is a crucial step in IE tasks. Existing models have achieved remarkable performance in the general domain Lin et al. (2020); Wang et al. (2021a); Zhang et al. (2023); Shen et al. (2023b). However, in the scientific domains, e.g., medical or chemical domains, these models usually struggle due to the extremely large quantity of concepts, the wide presence of multi-token entities, and the ambiguity in detecting entity boundaries.
@@
Figure 1: A test example in the chemical domain. The words marked with blue indicators are the chemical entities, and the words marked with red and orange indicators are biomedical entities. The entities in red are mis-labeled by a few-shot model as chemical entities.
@@
Large language models (LLMs) show an impressive performance on various NLP tasks such as question answering or text summarization OpenAI (2022). Models such as ChatGPT OpenAI (2022) can achieve outstanding results given just a few training examples Wang et al. (2022). However, Kandpal et al. (2023) recently report that the performance of these models is proportional to the number of relevant documents present in their pretraining corpus. Thus, one can expect that their performance fluctuates across domains. This is particularly expected to occur across certain scientific subjects, where the data may be scarce. Given the already existing challenges of the name tagging task in the scientific domain—mentioned earlier—this factor can further exacerbate the problem. For instance, in our early few-shot learning experiments, we observed that the results of ChatGPT in the chemical domain name tagging task are significantly worse than those in the general domain. [^1] We report this complementary experiment in Appendix A.
@@
In the present work, we employ transfer learning (Pan and Yang, 2010) to alleviate this problem. Transfer learning methods exploit the label data from one domain (called the source domain) to minimize the prediction error in another domain (namely the target domain). We particularly focus on a realistic setting, where given a large set of labeled data from the source domain and a small set of labeled data from the target domain, the goal is to develop a model for the target domain. [^2] In the literature, this setup is often called the semi-supervised transfer learning setting (Saito et al., 2019). There are more resources in the biomedical domain than in the chemical domain due to funding priorities and BioNLP workshops. Therefore, as a case study, we take the biomedical domain as the source and the chemical domain as the target.
@@
Figure 1 shows the challenges a name tagging model can face in the chemical domain. The model is trained in a few-shot learning setting. Thus, it is trained on labeled biomedical data, and then, further finetuned on a few labeled examples from the chemical domain. The task is to recognize only the chemical entities—ignoring other types of entities. Blue indicators represent chemical entities, while red and orange indicators denote biomedical entities. The model categorizes the entities marked with blue and red as chemical entities. The first observation is that, to an inexpert human, it is difficult to perform the task because the entities are highly domain-specific. The second observation is that the model wrongfully labels some entities from the source domain as the target entities. Therefore, engineers developing such a model face a dilemma: while not using the source data dramatically deteriorates performance, [^3] We empirically support this in the analysis section. simply pretraining with the source data increases the false positive rate. The third, and perhaps the most important, observation is that the examples from the source and target domains can co-occur in the same document. This problem setting contradicts the regular transfer learning setting, where the examples from the source and the target domains are fully disjoint (Ben-David et al., 2010). The latter property poses difficulties in the applicability of the traditional transfer learning methods in this setting.
@@
Our core idea is to train a name tagging model such that it is able to project the representations of the source and target entities into separate regions of the feature space. Such a model can potentially transfer knowledge from the source domain to the target domain by constructing a shared feature space between the two domains. Furthermore, it reduces the similarity between the representations of the entities in the two domains, and consequently, it can potentially minimize the number of source entities that are mislabeled as target entities. To achieve this, our model consists of a pretraining stage on the source data, and a finetuning stage on the target data.
@@
  • We propose a new pretraining algorithm for the name tagging task in the transfer learning setting. Our algorithm consists of two steps. First, extracting auxiliary information about the source entities through the event mentions they participate in. Second, proposing an entity grouping technique using the multi-similarity loss. Our methods have proven effective for the name tagging task in the target domain. Our study is carried out in the scientific domain, particularly from the biomedical data as the source domain to the chemical data as the target domain. As we stated before, this is a crucial and challenging real world problem. All of our claims, here and later, are only about this task.
  • We propose a finetuning algorithm for the name tagging task in the transfer learning setting. Our algorithm aims to project the target entities and the entities that may be potentially mislabeled into separate regions of the feature space. It consists of two steps. First, detecting the potentially out-of-domain entities by pseudo-labeling them. Second, obtaining dissimilar representations for the two sets of entities using the multi-similarity loss.
  • We conduct extensive experiments across twelve use cases, demonstrating that our method significantly outperforms the baseline models and shedding light on various aspects of our model.
Figure 2: Overview of proposed entity grouping and entity discrimination frameworks. Entity grouping on the source domain is shown in the upper part. Based on event annotations, a set of event embeddings is constructed under two paradigms. Afterward, pairwise auxiliary similarity scores are calculated according to argument embeddings. Extended multi-similarity loss concerning four types of similarities, combined with cross-entropy loss, are jointly learned during pretraining. Entity discrimination on the target domain is shown in the lower part. Pseudo labels are formed by the name tagging model pretrained in the source domain, and in contrast to annotated labels, a multi-similarity loss is injected into finetuning.
@@ 2.1 Named Entity Recognition
We view the name tagging task as a sequence labeling problem. We denote the training data by $D{=}\left\{(\mathbf{X}_{i},\mathbf{Y}_{i})\right\}_{i=1}^{n}$, where $n$ is the number of input passages (or texts). To train classifier parameter $\theta$, we minimize the loss as follows:
@@ 2.2 Transfer Learning
We focus on the name tagging task, and take the biomedical domain as the source and the chemical domain as the target. Note that the data distributions in the two domains are different. Therefore, a model solely trained on the source data is usually not as competitive as one trained on the target data. The baseline solution in this setting, direct transfer, is to pretrain the model on the labeled source data and finetune it on the labeled target data.
@@ 2.3 Multi-Similarity Loss
We enhance our name tagging model in the source domain by capturing the similarities between entity pairs. To achieve this, we employ an objective term called the multi-similarity (MS) contrastive loss proposed by Wang et al. (2019) for metric learning. MS can incorporate self-similarity, relative positive similarity, and relative negative similarity. Self-similarity depends on the properties of the data point itself, such as hardness, while negative and positive similarities are measured with respect to an anchor data point.
In the following sections, we use the final encoder hidden states of input tokens as entity representations. If an entity consists of multiple tokens, we take the average of their representations. Additionally, we use the cosine metric to measure the relative similarity between entity pairs.
@@
Second, a soft weight score is calculated for the extracted pairs to reflect their importance in the final loss term:
@@
where $n_{e}$ is the total number of the entities. Note that the values of $S_{\mathord{\color[rgb]{0.67,0.67,0.67}\bullet}}$ already contain the weights computed in Equations 3.
@@ 3 Proposed Method
Figure 2 illustrates our framework. It consists of two stages, pretraining on the source data, and then, finetuning on the target data. In the source domain, we employ external knowledge to construct an event feature space based on their arguments, and to calculate the auxiliary similarity scores between entities. Then, the similarity scores are used in the multi-similarity loss to shape the entity feature space. In the target domain, we aim to enhance the model ability to distinguish the target entities from the entities that are likely to be mislabeled, such as the entities that potentially belong to the source domain. To achieve this, we propose an algorithm to extract pseudo-labels and employ the multi-similarity loss for the second time.
@@ 3.1 Source Domain Pretraining
Using external knowledge, we enrich entity representations for source domain pretraining. Below, we propose two separate approaches to obtain the auxiliary embedding vectors, both exploiting event mentions that the entities appear in. Given the auxiliary vectors, next, we describe our method for calculating the auxiliary similarity scores. Finally, we provide an overview of the pretraining loss function, which incorporates the entity similarity scores and the auxiliary similarity scores.
Concatenation based Event Embedding.
Our first approach relies on an off-the-shelf token encoder pretrained on biomedical data, called SapBERT Liu et al. (2021a). Given an event mention, we encode its arguments using SapBERT, and then concatenate the resulting vectors to obtain the event representation. Note that in some cases an argument may be a nested event, or an event may have varying number of arguments. In those cases, we use vector averaging to compress the representations or padding to fill in the extra argument slots [^4] The details can be found in Appendix B.1..
@@
In our second approach, we use templates generated by a LLM. We begin by extracting all event types from the source domain. We then submit each type and its arguments to the LLM, using prompts to construct a template. A few examples of such templates are reported in Table 1, and a larger set of templates along our prompt instruction can be found in Appendix B.2. Then, for the event mentions in the source data, we complete their corresponding templates by replacing the placeholders with the actual arguments. The resulting passages are sent through an off-the-shelf sentence encoder to obtain the final representation vectors. In our experiments, we use ChatGPT OpenAI (2022) as the LLM, and use the S-PubMed-BERT Deka et al. (2022) as the sentence encoder.
In the result section, we individually evaluate each one of the methods for extracting the auxiliary embedding vectors.
Type
Template
Phosphorylation
Indicated by the given trigger <Trigger>, a specific molecule <Theme> is modified by the addition of a phosphate group at a particular site <Site>, facilitated by another molecule <Cause>.
Acetylation
Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes the addition of an acetyl group at a particular site <Site>, catalyzed by another molecule <Cause>.
Pathway
Indicated by the given trigger <Trigger>, involving one or more molecules <Participant> that collaborate to accomplish a specific biological function or response.
@@
where, as mentioned earlier, $\rho$ and $\tau$ are to maintain a balance between the primary and the auxiliary representations. Note that we use the information extracted from events to construct the auxiliary representations. However, additional source of information can be considered if it is present.
The pretraining objective function consists of the supervised name tagging term and the unsupervised RMS term, as follows:
@@ 3.2 Target Domain Finetuning
The source and target domains in our problem setting share the same context. In the same documents (or even sentences) that the entities from one domain appear, the entities from the other domain may be used too. This makes the recognition task particularly challenging in the target domain for two reasons: the training data in this domain is scarce, and the presence of entities from the source domain can potentially lead to a high false positive rate. Our core idea is that, while finetuning the model on the target data, we train the encoder such that it projects the target entities and the entities that potentially belong to the source domain into separate regions of the feature space. To implement this idea, we employ pseudo labeling along the multi-similarity loss–introduced in §2.3.
@@ 4.1 Experimental Setup
Dataset
# Train/ Valid/ Test
# Train/ Valid/ Test (All)
PC
260 / 89 / 175
ID
150 / 46 / 117
CG
300 / 100 / 200
CHEMDNER
81 / 72 / 2478
2916 / 2907 / 2478
BC5CDR
86 / 88 / 500
500 / 500 / 500
DrugProt
93 / 88 / 600
597 / 597 / 600
@@
Target Tasks
CHEMDNER BC5CDR DrugProt
Metrics
Precision
Recall
F1
Precision
Recall
F1
Precision
Recall
F1
Target Only
42.73
51.69
46.77
72.44
85.86
78.51
63.80
67.42
65.49
Direct Transfer
44.11
48.60
46.18
71.92
86.54
78.55
66.36
70.13
68.17
EG(concat)-Only
43.63
51.06
47.03
72.10
87.12
78.89
68.49
66.54
67.49
EG(sentEnc)-Only
43.09
51.60
46.91
73.96
86.56
79.76
66.02
69.97
67.93
ED-Only
45.71
50.13
47.78
75.01
87.80
80.88
70.15 71.79 70.96
EG(concat)+ED
43.83
52.17
47.64
73.70
85.50
79.15
68.01
69.06
68.52
EG(sentEnc)+ED
45.28
51.68
48.26 76.06
86.58
80.95
67.16
69.04
68.08
@@
Datasets. As the source datasets, we use three benchmarks: Pathway Curation (PC), Cancer Genetics (CG), and Infectious Diseases (ID). The first two datasets were released by BioNLP Shared Task 2013 Nédellec et al. (2013), and the third one was released by BioNLP Shared Task 2011 Pyysalo et al. (2011). All three datasets have the same format. We aggregate them to create a fourth dataset called the biomedical multi-task dataset. As the target datasets, we use three benchmarks: CHEMDNER Krallinger et al. (2015), BC5CDR Kim et al. (2015), and DrugPort Miranda et al. (2021).
Each dataset contains extra annotations unrelated to its domain. For instance, the CG dataset has annotations not relevant to biomedicine.
We pre-process all the datasets by removing such annotations.
For few-shot experiments, we down-sample the training and validation sets of target datasets to sizes randomly chosen between 70 and 100. Table 2 summarizes the dataset statistics.
@@
We compare our method with two baselines. Target Only: a model finetuned on the labeled target data.
Direct Transfer: a model pretrained on the labeled source data and then finetuned on the labeled target data.
@@
We use BERT (base-uncased) Devlin et al. (2019) as the backbone for all the models. To train the model, we only update the parameters of the adapter layers Houlsby et al. (2019) and freeze the rest. We iteratively select each source and target dataset pair as the training and evaluation benchmarks. All the experiments are repeated three times. Following Nakayama (2018), we report average macro Precision, Recall, and F1 scores [^5] Complementary information about training and tuning can be found in Appendix C..

### v2

# Named Entity Recognition Under Domain Shift via Metric Learning for Life Sciences

@@ Abstract
Named entity recognition is a key component of Information Extraction (IE), particularly in scientific domains such as biomedicine and chemistry, where large language models (LLMs), e.g., ChatGPT, fall short. We investigate the applicability of transfer learning for enhancing a named entity recognition model trained in the biomedical domain (the source domain) to be used in the chemical domain (the target domain). A common practice for training such a model in a few-shot learning setting is to pretrain the model on the labeled source data, and then, to finetune it on a hand-full of labeled target examples. In our experiments, we observed that such a model is prone to mislabeling the source entities, which can often appear in the text, as the target entities. To alleviate this problem, we propose a model to transfer the knowledge from the source domain to the target domain, but, at the same time, to project the source entities and target entities into separate regions of the feature space. This diminishes the risk of mislabeling the source entities as the target entities. Our model consists of two stages: 1) entity grouping in the source domain, which incorporates knowledge from annotated events to establish relations between entities, and 2) entity discrimination in the target domain, which relies on pseudo labeling and contrastive learning to enhance discrimination between the entities in the two domains. We conduct our extensive experiments across three source and three target datasets, demonstrating that our method outperforms the baselines by up to 5% absolute value [^1] Code, data, and resources are publicly available for research purposes: https://github.com/Lhtie/Bio-Domain-Transfer..
Named Entity Recognition Under Domain Shift via Metric Learning
@@
Hongyi Liu1,  Qingyun Wang2,  Payam Karisani2,  Heng Ji2 1 Shanghai Jiao Tong University, 2 University of Illinois at Urbana-Champaign liu.hong.yi@sjtu.edu.cn, {qingyun4,karisani,hengji}@illinois.edu
@@ 1 Introduction
Named entity recognition is a crucial step in IE tasks. Existing models have achieved remarkable performance in the general domain Lin et al. (2020); Wang et al. (2021b); Zhang et al. (2023); Shen et al. (2023b). However, in the scientific domains, e.g., medical or chemical domains, these models usually struggle due to the extremely large quantity of concepts, the wide presence of multi-token entities, and the ambiguity in detecting entity boundaries.
@@
Figure 1: A test example in the chemical domain. The words marked with blue indicators are chemical entities, and the words marked with red and orange indicators are biomedical entities. The entities in red are mislabeled by a few-shot model as chemical entities.
@@
Large language models (LLMs) show an impressive performance on various NLP tasks such as question answering or text summarization OpenAI (2022). Models such as ChatGPT OpenAI (2022) can achieve outstanding results given just a few training examples Wang et al. (2022). However, Kandpal et al. (2023) recently report that the performance of these models is proportional to the number of relevant documents present in their pretraining corpus. Thus, one can expect that their performance fluctuates across domains. This is particularly expected to occur across certain scientific subjects, where the data may be scarce. Given the already existing challenges of the named entity recognition task in the scientific domain—mentioned earlier—this factor can further exacerbate the problem. For instance, in our early few-shot learning experiments, we observed that the results of ChatGPT in the chemical domain named entity recognition task are significantly worse than those in the general domain. [^2] We report this complementary experiment in Appendix A.
@@
In the present work, we employ transfer learning (Pan and Yang, 2010) to alleviate this problem. Transfer learning methods exploit the label data from one domain (called the source domain) to minimize the prediction error in another domain (namely the target domain). We particularly focus on a realistic setting, where given a large set of labeled data from the source domain and a small set of labeled data from the target domain, the goal is to develop a model for the target domain [^3] In the literature, this setup is often called the semi-supervised transfer learning setting (Saito et al., 2019).. There are more resources in the biomedical domain than in the chemical domain due to funding priorities and BioNLP workshops. Therefore, as a case study, we take the biomedical domain as the source and the chemical domain as the target.
@@
Figure 1 shows the challenges a named entity recognition model can face in the chemical domain. The model is trained in a few-shot learning setting. Thus, it is trained on labeled biomedical data, and then, further finetuned on a few labeled examples from the chemical domain. The task is to recognize only the chemical entities—ignoring other types of entities. Blue indicators represent chemical entities, while red and orange indicators denote biomedical entities. The model categorizes the entities marked with blue and red as chemical entities. The first observation is that, to an inexpert human, it is difficult to perform the task because the entities are highly domain-specific. The second observation is that the model wrongfully labels some entities from the source domain as the target entities. Therefore, engineers developing such a model face a dilemma: while not using the source data dramatically deteriorates performance, [^4] We empirically support this in the analysis section. simply pretraining with the source data increases the false positive rate. The third, and perhaps the most important, observation is that the examples from the source and target domains can co-occur in the same document. This problem setting contradicts the regular transfer learning setting, where the examples from the source and the target domains are fully disjoint (Ben-David et al., 2010). The latter property poses difficulties in the applicability of the traditional transfer learning methods in this setting.
@@
Our core idea is to train a named entity recognition model such that it is able to project the representations of the source and target entities into separate regions of the feature space. Such a model can potentially transfer knowledge from the source domain to the target domain by constructing a shared feature space between the two domains. Furthermore, it reduces the similarity between the representations of the entities in the two domains, and consequently, it can potentially minimize the number of source entities that are mislabeled as target entities. To achieve this, our model consists of a pretraining stage on the source data, and a finetuning stage on the target data.
@@
  • We propose a new pretraining algorithm for the named entity recognition task in the transfer learning setting. Our algorithm involves two steps: first, extracting auxiliary information about the source entities through the event mentions they participate in; and second, proposing an entity grouping technique using the multi-similarity loss. Our methods have proven effective for the named entity recognition task in the target domain. Our study is carried out in the scientific domain, particularly from the biomedical data as the source domain to the chemical data as the target domain. This is a crucial and challenging real-world problem. All of our claims, here and later, are only about this task.
  • We propose a finetuning algorithm, which aims to project the target entities and the entities that may be potentially mislabeled into separate regions of the feature space. It comprises two steps: first, detecting the potentially out-of-domain entities by pseudo-labeling them; and second, obtaining dissimilar representations for the two sets of entities using the multi-similarity loss.
  • We conduct extensive experiments across twelve cases, showing that our method significantly outperforms the baselines and shedding light on various aspects of our model.
Figure 2: Overview of proposed entity grouping and entity discrimination frameworks. Entity grouping on the source domain is shown in the upper part. Based on event annotations, a set of event embeddings is constructed under two paradigms. Afterward, pairwise auxiliary similarity scores are calculated according to argument embeddings. Extended multi-similarity loss concerning four types of similarities, combined with cross-entropy loss, are jointly learned during pretraining. Entity discrimination on the target domain is shown in the lower part. Pseudo labels are formed by the named entity recognition model pretrained in the source domain, and in contrast to annotated labels, a multi-similarity loss is injected into finetuning.
@@ 2.1 Named Entity Recognition
We view the named entity recognition task as a sequence labeling problem. We denote the training data by $D{=}\left\{(\mathbf{X}_{i},\mathbf{Y}_{i})\right\}_{i=1}^{n}$, where $n$ is the number of input passages (or texts). To train a classifier $f$ with parameter $\theta$, we minimize the loss as follows:
@@ 2.2 Transfer Learning
We focus on the named entity recognition task, and take the biomedical domain as the source and the chemical domain as the target. Note that the data distributions in the two domains are different. Therefore, a model solely trained on the source data is usually not as competitive as one trained on the target data. The baseline solution in this setting, direct transfer, is to pretrain the model on the labeled source data and finetune it on the labeled target data.
@@ 2.3 Multi-Similarity Loss
We enhance our named entity recognition model in the source domain by capturing the similarities between entity pairs. To achieve this, we employ an objective term called the multi-similarity (MS) contrastive loss proposed by Wang et al. (2019) for metric learning. MS can incorporate self-similarity, relative positive similarity, and relative negative similarity. Self-similarity depends on the properties of the data point itself, such as hardness, while negative and positive similarities are measured with respect to an anchor data point.
In the following sections, we use the final encoder hidden states of input tokens as entity representations. If an entity consists of multiple tokens, we take the average of their representations. Additionally, we denote the relative similarity between entity pairs as $S_{\mathord{\color[rgb]{0.67,0.67,0.67}\bullet}}$, using cosine similarity.
@@
Second, we calculate a soft weight score for the extracted pairs to reflect their importance:
@@
where $n_{e}$ is the total number of the entities. Note that the values of $S_{\mathord{\color[rgb]{0.67,0.67,0.67}\bullet}}$ already contain the weights computed in Equation 3.
@@ 3 Proposed Method
Figure 2 illustrates our framework. It consists of two stages, pretraining on the source data, and then, finetuning on the target data. In the source domain, we employ external knowledge to construct an event feature space based on their arguments, and to calculate the auxiliary similarity scores between entities. Then, the similarity scores are used in the multi-similarity loss to shape the entity feature space. In the target domain, we aim to enhance the model’s ability to distinguish the target entities from the entities that are likely to be mislabeled, such as the entities that potentially belong to the source domain. To achieve this, we propose an algorithm to extract pseudo-labels and employ the multi-similarity loss for the second time.
@@ 3.1 Source Domain Pretraining
Using external knowledge, we enrich entity representations for source domain pretraining. Since the source domain consists of a set of various sub-domains (or sub-topics), discovering these sub-domains in the source domain facilitates the subsequent process of domain transfer Hoffman et al. (2012). Specifically, by detecting these sub-domains and grouping the data, we enable the contrastive loss in the next step to consider each one individually and transform them accordingly. This approach avoids the oversimplification of treating the entire source domain as a single cluster. Below, we propose two separate approaches to obtain the auxiliary embedding vectors, both exploiting event mentions that the entities appear in. Given the auxiliary vectors, we describe our method for calculating the auxiliary similarity scores. Finally, we provide an overview of the pretraining loss function, which incorporates the entity similarity scores and the auxiliary similarity scores.
Concatenation-based Event Embedding.
Our first approach relies on an off-the-shelf token encoder pretrained on biomedical data, called SapBERT Liu et al. (2021a). Given an event mention, we encode its arguments using SapBERT, and then concatenate the resulting vectors to obtain the event representation. Note that in some cases, an argument may be a nested event, or an event may have a varying number of arguments. In those cases, we use vector averaging to compress the representations or padding to fill in the extra argument slots [^5] The details can be found in Appendix C.1..
@@
In our second approach, we use templates generated by a LLM. We begin by extracting all event types from the source domain. We then submit each type and its arguments to the LLM, using prompts to construct a template. A few examples of such templates are reported in Table 1, and a larger set of templates along our prompt instruction can be found in Appendix C.2. Then, for the event mentions in the source data, we complete their corresponding templates by replacing the placeholders with the actual arguments. The resulting passages are sent through an off-the-shelf sentence encoder to obtain the final representation vectors. In our experiments, we use ChatGPT OpenAI (2022) as the LLM, and use the S-PubMed-BERT Deka et al. (2022) as the sentence encoder. In § 4, we individually evaluate each one of the methods for extracting the auxiliary embedding vectors.
Type Template
Phosphorylation Indicated by the given trigger <Trigger>, a specific molecule <Theme> is modified by the addition of a phosphate group at a particular site <Site>, facilitated by another molecule <Cause>.
Acetylation Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes the addition of an acetyl group at a particular site <Site>, catalyzed by another molecule <Cause>.
Pathway Indicated by the given trigger <Trigger>, involving one or more molecules <Participant> that collaborate to accomplish a specific biological function or response.
@@
where, as mentioned earlier, $\rho$ and $\tau$ are to maintain a balance between the primary and the auxiliary representations. Note that we use the information extracted from events to construct the auxiliary representations. However, additional sources of information can be considered if it is present.
The pretraining objective function consists of the supervised named entity recognition term and the unsupervised RMS term, as follows:
@@ 3.2 Target Domain Finetuning
The source and target domains in our problem setting share the same context. In the same documents (or even sentences) that the entities from one domain appear, the entities from the other domain may be used, too. This makes the recognition task particularly challenging in the target domain for two reasons: the training data in this domain is scarce, and the presence of entities from the source domain can potentially lead to a high false positive rate. Our core idea is that, while finetuning the model on the target data, we train the encoder such that it projects the target entities and the entities that potentially belong to the source domain into separate regions of the feature space. To implement this idea, we employ pseudo labeling along the multi-similarity loss–introduced in §2.3.
@@ 4.1 Experimental Setup
Dataset # Train/ Valid/ Test # Train/ Valid/ Test (All)
PC - 260 / 89 / 175
ID - 150 / 46 / 117
CG - 300 / 100 / 200
CHEMDNER 81 / 72 / 2478 2916 / 2907 / 2478
BC5CDR 86 / 88 / 500 500 / 500 / 500
DrugProt 93 / 88 / 600 597 / 597 / 600
@@
Target Tasks CHEMDNER BC5CDR DrugProt
Metrics Precision Recall F1 Precision Recall F1 Precision Recall F1
Target Only 42.73 51.69 46.77 72.44 85.86 78.51 63.80 67.42 65.49
Direct Transfer 44.11 48.60 46.18 71.92 86.54 78.55 66.36 70.13 68.17
EG(concat) 43.63$\pm$1.7 51.06$\pm$0.1 47.03$\pm$1.0 72.10$\pm$1.3 87.12$\pm$0.5 78.89$\pm$0.6 68.49$\pm$0.9 66.54$\pm$0.9 67.49$\pm$0.2
EG(sentEnc) 43.09$\pm$3.7 51.60$\pm$3.3 46.91$\pm$2.9 73.96$\pm$1.1 86.56$\pm$0.8 79.76$\pm$0.3 66.02$\pm$0.5 69.97$\pm$0.8 67.93$\pm$0.3
ED 45.71$\pm$2.3 50.13$\pm$1.2 47.78$\pm$1.3 75.01$\pm$2.0 87.80$\pm$0.5 80.88$\pm$0.9 70.15$\pm$0.3 71.79$\pm$0.3 70.96$\pm$0.1
EG(concat)+ED 43.83$\pm$0.6 52.17$\pm$0.9 47.64$\pm$0.7 73.70$\pm$0.8 85.50$\pm$0.9 79.15$\pm$0.1 68.01$\pm$0.2 69.06$\pm$1.3 68.52$\pm$0.6
EG(sentEnc)+ED 45.28$\pm$0.4 51.68$\pm$1.8 48.26$\pm$0.9 76.06$\pm$1.8 86.58$\pm$1.3 80.95$\pm$0.5 67.16$\pm$0.8 69.04$\pm$1.0 68.08$\pm$0.9
@@
Datasets.
As the source datasets, we use three benchmarks: Pathway Curation (PC), Cancer Genetics (CG), and Infectious Diseases (ID). The first two datasets were released by BioNLP Shared Task 2013 Nédellec et al. (2013), and the third one was released by BioNLP Shared Task 2011 Pyysalo et al. (2011). All three datasets have the same format. We aggregate them to create a fourth dataset called the biomedical multi-task dataset. As the target datasets, we use three benchmarks: CHEMDNER Krallinger et al. (2015), BC5CDR Kim et al. (2015), and DrugPort Miranda et al. (2021).
Each dataset contains extra annotations unrelated to its domain. For instance, the CG dataset has annotations not relevant to biomedicine. We pre-process all the datasets by removing such annotations. For few-shot experiments, we down-sample the training and validation sets of target datasets to sizes randomly chosen between 70 and 100. Table 2 summarizes the dataset statistics.
@@
We compare our method with two baselines. Target Only: a model finetuned on the labeled target data. Direct Transfer: a model pretrained on the labeled source data and then finetuned on the labeled target data.
@@
We use BERT (bert-base-uncased) Devlin et al. (2019) as the backbone for all the models. To train the model, we update the parameters of the adapter layers Houlsby et al. (2019) and freeze the rest, due to limited computational resources. We iteratively select each source and target dataset pair as the training and evaluation benchmarks. All the experiments are repeated three times. Following Nakayama (2018), we report average macro Precision, Recall, and F1 scores [^6] Training and tuning details can be found in Appendix D..